TWiki home Point 1 Point 2 Point 3 Point 4 Point 5 Point 6 Point 7 Point 8 Minutes

Back to overview

YOU ARE VIEWING BlogEntryPoint8x2007x05x24x18x26

Powering tests in Sector 7-8: activity of Thursday 24/05.

The day was fully devoted to perform tests on Q4, which had experienced an unexpected quench at 300 A on Monday (after ramping at 600 A).

The following tests have been executed on Q4 (MANUAL execution, due to change in procedure):
  • ramp to 600 A with 5 A/s and quench driven by QPS, generating an artificial quench signal from one current lead detector; the test was successful;
  • ramp of B1 power converter to nominal current (3610 A) with 10 A/s ramp rate and simultaneous ramp of B2 power converter to 2000 A with nearly 5 A/s (unbalanced configuration):
    • when B1 current reached 1025 A, its reference current went anormally down to 0;
    • later analysis showed that a fault was sent from the converter and a consequent FPA was started;
    • PO attempt to get control of B2 power converter, to ramp it down contemporarily to 0, provoked a sudden voltage peak in B1, which initiated a quench;
  • at 11.48 (according to PIC supervision) a power cut occured on the general services of UA83, which eventually provoked the loss of controls on cryogenics and lead heating system;
  • after waiting for the good cryo conditions to be recovered, and in agreement with PO, QPS and MPP experts, we decided to ramp again Q4 with the same unbalanced cycle:
    • B1 at 3610 A with 10 A/s ramp rate and B2 to 2000 A with a nearly 5 A/s ramp rate;the test was successful and NO quench was activated;
    • while ramping B2 power converter to the same current as B1 (the idea was to stay on plateau for half an hour and then provoke a quench), and with a current in B2 of 2900 A, a quench generated; preliminary analysis indicated that it was started in B1, possibly due to electromagnetic coupling (this configuration has never been tested before).
  • end of the tests due to long recovery time.

Proposal for the activity of tomorrow.

  • Continue powering Q4 and D2.

  • Continue installation of plastic bags on 600 A current leads.
  • Lower the temperature on the 13 kA leads to check the absence of condensation with AT/MEL copper laminations.

  • ElQA on the selected circuits: list will be provided during the morning.

Edit | Attach | Watch | Print version | History: r2 < r1 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r2 - 2007-05-25 - MirkoPojer
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    HCC All webs login